OpenStudy (anonymous):

Derive the quadratic function if the vertex has coordinates (4,4) containing the point (o,-2)

OpenStudy (amistre64):

y = m(x-4)^2+4 m = (y-4)/(x-4)^2 m = (-2-4)/(0-4)^2

OpenStudy (anonymous):

is it the final answer??